The difference between two numbers is 3. If the sum of twice the larger and the smaller is 48, find the two numbers. this is how I tried to solve it.
2n + n - 3 =48 n = 15 Thanks in advance for your help.

Answer by mananth(16949)

The difference between two numbers is 3. If the sum of twice the larger and the smaller is 48, find the two numbers. this is how I tried to solve it.
2n + n - 3 =48 n = 15 Thanks in advance for your help.
let one number be x
the other number will x-3
2x+x-3=48
3x=48+3 ( sign changes when you shift it to the other side )
3x=51
x= 51/3
x=17
the other number will be 14
2*17+14=48
ananth